OpenStudy (anonymous):

the sum of two numbers is 40. the smaller number is 22 less than the larger number. what are the numbers?

OpenStudy (amistre64):

(20+n) + (20-n) = 40 (20-n) = (20+n) - 22 solve for n to determine the number .... maybe